An inelastic collision occurs in one dimension, in which a 10 kg block traveling at

5 m/s collides with a 5 kg block traveling at 3 m/s in the same direction, and they

stick together. What are the velocities of the blocks immediately after the collision?


1
Expert's answer
2021-09-12T19:15:30-0400

Solution


From the law of conservation of momentum


"m_1 V_{1i} + m_2. V_{2i} = m_1V_1 + m_2V_2"


where vii- velocities before collision are v1i=5 m/s Particles of "m_1=10 kg"

"V_{2i}-" velocities before collision are "v_{2i}=3 m\/s" Particles of "m_2=5 kg"

"V_i -" velocities after collision Particles of "m_1=10 kg"

"V_2 -" velocities after collision Particles of "m_2=5 kg"


Energy conservation law


"m_1 V_{1i}^2\/2 + m_2. V_{2i}^2\/2 = m_1V_1^2\/2 + m_2V_2^2\/2"


from here:

"v_1=\frac{v_{1i}(m_1-m_2)}{m_1+m_2}+\frac{v_{2i}(2)(m_2)}{m_1+m_2}=\frac{5(10-5)}{10+5}+\frac{3\times2\times5}{10+5}=3.66m/s"


"v_2=\frac{v_{1i}(2)(m_1)}{m_1+m_2}+\frac{v_{2i}(m_2-m_1)}{m_1+m_2}=\frac{5\times2\times10}{10+5}+\frac{3(5-10)}{10+5}+=5.66m/s"


ANSWER:

"v_1=3.66m\/s"


"v_2=5.66m\/s"